OpenStudy (cicilybailey14):

Fan+Medal! Write an equation of the direct variation that includes the given point: (1, 5) y = 5x + 1 y = 5x - 1 y = 1/5x y = 5x

OpenStudy (ganpat):

direct variation... y is directly proportional to x = (y=Kx)| Where K is proportional constant and points are given in the form (x,y)

OpenStudy (michele_laino):

note that we have, by definition: y=k*x so substituting your point, we can write: 5=k*1 so k=5 so y=5*x

OpenStudy (cicilybailey14):

@Michele_Laino thanks!

OpenStudy (michele_laino):

@Cicilybailey14 thanks!